$2,000.00 invested at 6% compounded quarterly; find the account balance at the end of the second period.
I have tried to multiply .06(2000) then add it to 2000, but it doesn't work. Or doing that then multiplying that number by 2, and that doesn't work. I know the correct answer is $2,060.45 but I need to know how to do it.

Answer by DrBeeee(684)

The 6% is for one year. The quarterly rate is one fourth of this or 1.5% or 0.015 as a decimal. Then your formula for compound interest on a quarterly basis is
(1) B = P*(1+.015)^n, where n is the number of quarters
In your example P = $2000 and n = 2 (the second period). Then you have
(2) B = 2000*(1.015)^2 or
(3) B = 2000*(1.030225) or
(4) B = 2060.45
Answer: At the end of the second period the balance of the account is $2,060.45
Comment: Always remember that the interest rate is (usually) given as the annual rate or by the year. If you want the monthly rate divide by 12. This (monthly) is used for mortgages which are paid monthly.
I hope this helped.